OpenStudy (anonymous):

Solve this equations for x: 8(x + 1) = 8(x – 1) + 2x A. x=1/2 B. x=-1 C. x=0 D. x=8

OpenStudy (shaik0124):

8x+8=8x-8+2x 8x+8-8x=8x-8+2x-8x 8=-8+2x 8+8=-8+8+2x 16=2x x=16/2 x=8

OpenStudy (shaik0124):

@venom936

OpenStudy (anonymous):

What is the domain of x in this expression? y+2/x+3 + √16x